OpenStudy (anonymous):

Find the distance between the points (3, 0) and (-1, 5).

OpenStudy (anonymous):

Try using the distance formula

OpenStudy (anonymous):

\[\sqrt{(-1-3)^2+(5-3)^2}\]

OpenStudy (anonymous):

\[\sqrt{10+34}\]

OpenStudy (anonymous):

Then I got to \[\sqrt{44}\]

OpenStudy (anonymous):

But how do you simplify that?

OpenStudy (cggurumanjunath):

\[\sqrt{(-1-3)^2+(5-0)^2}\]

OpenStudy (cggurumanjunath):

\[\sqrt{(-1-3)^2+(5-0)^2}\] = sqrt (16+25)

OpenStudy (cggurumanjunath):

sqrt (41)

OpenStudy (cggurumanjunath):

sqrt(41) is the final answer @ilmc1

OpenStudy (anonymous):

Thank you again :)

OpenStudy (cggurumanjunath):

:)